OpenStudy (anonymous):

Two planes leave an airport at noon. If the eastbound plane flies at 560 mph and the westbound plane flies at 500 mph, at what time will they be 2,000 miles apart? Round your answer to the nearest minute.

OpenStudy (anonymous):

560x+500x=2,000 1060x=2,000 x=2000/1060 x=1.88679 hours or they'll be 2,000 miles

OpenStudy (anonymous):

damn, i spent my time doing calculus